what 2 digit number is equal to twice the product of its digits?
Answer by RAY100(1637)

XY =2xy
.
x=3,,y=6
.
check 36 = 2(3*6) =2*18=36
.
However, an exact solution is difficult.
.
y,, must be an even digit,,as 2*xy will end in an even number
.
y cannot = 0, (unless we will satisfy for the trivial 0=0),,, as we multiply by y in 2xy
.
Is there any other info, to help define an exact solution ?